Volume Imaging Lidar Status at Lake-ICE

This page contains information on the current status of the Volume Imaging Lidar.
Return to the Index. Return to Current Operations Menu.
Check out details of VIL Operations in Lake-ICE here.

Lake-Induced Convection Experiment, Sheboygan, WI

  • 1/19/98: 3 pm CST: crew decided to return to Madison. Operation of laser may cause further damage to other optics.

  • 1/19/98: 1:25 pm CST: VIL collected data today from 12:24-19:06 Z. Laser power dropped from 350 mJ/pulse at start of today to 315 mJ/pulse at 19:06 Z. (Typical laser power is 400 mJ/pulse.) Inspection of laser revealed tiny crack on the side of one of the laser rods and a blemish on one of the internal lens. Cleaning the lens did not help--if fact, laser power decreased to about 250 mJ/pulse.

    During observations this morning, we saw growth of the mixed layer from 200 m AGL to about 500 m AGL over several hours. The height of the mixed layer is fairly constant spatially, i.e. NOT deepening with distance offshore. Very good data today-especially later in the morning and throughout the mid-day.

  • 1/19/98: VIL collecting data since 12:24 Z (6:24 am CST). Overcast skies with flurries during early morning. Complex ABL.

  • 1/18/98: 18:10 Z (12:10 pm CST) Shutdown VIL for the remainder of today. Approximately 6 hours of data collected this morning.

  • 1/18/98: 12 Z (6 am CST) Arrived at VIL site at 3:15 am CST. Trouble with laser and computer delayed start of data collection until 4:49 am CST. ABL is about 200 m deep and there is a hint of a TIBL deepening with increasing distance offshore in the VIL data. Winds are offshore (from 280 degrees) at 13 knots. Air temperature is 19.6 F (dew point 15.3 F).

  • 1/17/98: VIL crew leaving Madison Saturday evening for Sheboygan. Research flights planned for 4:45 am EST (3:45 am CST) departure in Ann Arbor. Hope to have the VIL up and running by approx 3 am CST. SDM.

  • 1/16/98: No evidence of either the Grey or Orange buoy was seen during the recovery operations on 1/15/98. The search started at approximately 2:30 p.m. (CST) and was terminated as 5:30 p.m. due to darkness. The sky was overcast, visibility was about 5 miles, winds were from the north at 10 knts and the waves were 2 - 3 feet. Four people were onboard the boat and continously looking.

    During the search the boat passed over the last known Argos position at least six times. The first hour of the search was spent within 1/4 mile of the last known position. During the second and third hours the search included a four mile square area centered on the last known postion of the Grey Niiler buoy. Jim Boyle

  • 1/14/98 3:00 pm CST. 1/15 is scheduled to be a down day for the aircraft. Weather in Sheboygan took a turn for the worse today with easterly flow and snow. Visibility very poor and sea rough. Lidar crew leaving Sheboygan at 3:15 PM and returning to Madison until next IOP.

  • 1/14/98 9:00 am CST. The optical damage that occurred on 1/11/98 was repaired before dawn on 1/13/98--just in time for data collection on Tuesday. Unfortunately, other mechanical problems (laser water chiller and beam-steering-unit became unreliable as the temperatures dropped to -5.6 F by 9 am CST) prevented us from collecting continuous data during a very good cold air outbreak. We managed to overcome the chiller and BSU problems enough to collect about 3 hours of data within the time span ranging from 12:47 to 18:37 Z on 1/13/98. Lidar crew rested Tuesday afternoon and evening.

    Weather conditions in Sheboygan on the morning of 1/14/98 (Wednesday) are overcast with snow (visibility about 3 or 4 km). When we arrived at the lidar van at 8 am CST we noticed westerly (offshore) flow at 4 m/s which changed to an ESE flow (off the lake) at 8 m/s by 8:45 am CST. By 9:30 am snow had stopped and visibility was much higher. Clouds appear to be stratocumulus now. Flow is onshore. (SDM)

  • 1/11/98 11:00 am CST. Collected very good RHI volume scans of thermal internal boundary layer and shallow offshore BL during three different phases 5:51-6:16; 7:47-8:30; and 9:29-10:23 am CST. Had problems with laser cooling system throughout morning which we managed to fix. Unfortunately, much more serious optical/laser problems (burned lens, mirrors, and destroyed a thin-film beam-splitter) occurred which completely stopped VIL operations for remainder of today and Monday. We hope to be operational again Tuesday afternoon. (We will try to obtain replacement parts before Tuesday morning, but can't promise.)

    Jim Boyle says the surface heat flux sensor is reporting data, but he thinks the values are too small. (Possible ice formation on the sensor?) (SDM)

  • 1/10/98 2:13 pm CST. We acquired lidar data from 12:37 UT to 18:47 UT. Good lidar backscatter structure at altitudes below 800 m should allow derivation of wind profiles as a function of distance from shore. Wide angle scans just above the lake surface provided images of the steam fog structure over a wide area.

    The heat flux buoy crew returned from a cruise on the lake. They were unsucessful in finding the buoy which they deployed on the 6th. Fisherman told them that the storm on Thursday had dragged their nets and it was likely that the heat flux buoy had also moved. The search for the first buoy will have to resume after the steam fog clears. A second heat flux buoy was deployed. This buoy is located at:

    43 deg 42.88 min N
    87 deg 39.73 min W
    The first data stream from the buoy should be received from the ARGOS station late this afternoon.

    The lidar is operational. Some system testing is planned for this afternoon. Ralph Kuehn and Shane Mayor will remain in Sheboygan with plans to operate the lidar tomorrow morning. Ed Eloranta will leave this afternoon and be out of town all of next week. Ralph and Shane will continue lidar operations during the week.

  • 1/9/98 1:40 pm CST. Part of the lidar crew is in route to the lidar to make another attempt to establish the PPP network connection to our Madison computers. The rest of the crew will follow this afternoon for possible short sequence operations this evening and extended operations beginning tomorrow morning.

    A second crew supporting the surface heat flux buoy is also in route to Sheboygan with plans to deploy a second buoy and to investigate why the first buoy stopped transmitting after only 8 hours of operation. This operation depends on finding a fishing boat captian willing to visit the site. They hope to get on the water this afternoon or early tomorrow.

    For those unfamiliar with the buoy, It measures heat conducted through the surface layer of the water. This provides a measure of the total surface heat flux including sensible, latent and radiative contributions. The buoy also provides a water temperature at a depth of approximately 1 cm and an air temperature at an altitude of approximately 1 cm. The air temperature measurement is difficult to interpret because the sensor is not always dry.

  • 1/8/98 2:07 pm CST. Several new mpeg loops showing lidar data obtained from Sheboygan have been added to our web page. Click on VIL observations. You will need to have an mpeg viewer installed in your browser to see these. The default mpeg viewer supplied with Microsoft Explorer has limitations which cause it to fail on some of these animations. We use the VMPEG viewer. A web search on Altavista will provide the URL from which the VMPEG code can be downloaded.

  • 1/7/98 8:00 am CST. Yesterday we deployed the heat flux buoy. We also installed a temperature controller on the lidar detector package to prevent temperature dependent drift of the preamp offset voltage. The lidar is ready for operation with the crew standing by in Madison. Ed Eloranta

    At 2:45 p.m. (CST) on 1/6/98 sensor float 005 was attached to the Orange Niiler buoy and moored in approximately 65 feet of water in Lake Michigan near Sheboygan. The coordinates from the Garmin GPS are:
    43d 42m 63.1s N
    87d 39m 36.1s W
    Data should be available after 6:00 a.m. Wednesday (1/7/98). Jim Boyle

  • 1/6/98 7:00 am CST. Jim Boyle and Ralph Kuehn plan to deploy a heat flux sensing buoy approximately 2 miles offshore from the lidar van this afternoon. The buoy will provide measurements of surface heat flux and temperature. Continuous measurements will be acquired via an ARGOS satellite data link. Battery life should be adequate for at least ten days of operation. The proposed location is 43 deg 42 min 49.2 sec N, 87 deg 39 min 44.6 sec W. We will provide exact coordinates for the buoy after installation. Aircraft measurements of the surface heat flux over the buoy would be useful if these can be arranged. The lidar is operational and the crew is standing by in Madison waiting for suitable weather.

  • 12/24/97 7:30 am CST No operations anticipated before Jan 2. We have experienced difficulties with the PPP network connection from the lidar van to our lab computers. This has no direct impact on data acquisition but it has prevented real time transfer of lidar images to our web page. The lidar is working well and we have recorded some very interesting data showing the land breeze front and the acceleration of the boundary layer flow as westerly winds leave the shore.

  • 12/19/97 5:15 pm CST Lidar crew leaving for Sheboygan in anticipation of operations tomorrow morning. Ed Eloranta

  • 12/15/97 1:00 pm CST Lidar obtained an interesting data set yesterday. The system is operational and the crew is standing by in Madison waiting for suitable weather. Ed Eloranta

  • 12/10/97 9:30 am CST Lidar crew is standing by in Madison waiting for suitable weather. Jim and Patrick at site today checking alignment and installing detector package temperature control. Ed Eloranta

  • 12/06/97 9:17 am CST UW lidar is not operating today- crew taking a well deserved break after a week of very long days in the lidar van. Ed Eloranta

  • 12/05/97 10:00 am CST UW lidar began operation at 15:32 GMT. Very complex aerosol and cloud structure which vaires strongly with direction. Attenuation in very light snow limits measurement range to about 7 km. The VIL is operational. Ed Eloranta

  • 2/04/97 10:01 am CST The telephone lines are now installed in lidar van (920-803-0605). Energy moniter calibrated. Progress was slowed by broken drive coupling on water pump. This is now repaired and we are begining log amp calibration. We believe that the system is marginally operational but have not tested full operation. Ed Eloranta

  • 12/03/97 9:30 am CST The beam steering unit is installed and electrical power is connected. System testing and alignment has begun. Problems caused by optical damage to one of the beam delivery mirrors have been corrected by moving the mirror to use an undamaged area. Quite a bit of work remains before the system is operational. Rain and snow are slowing the alignment of optics in the beam steering unit. We still hope to have to lidar working late today. Ed Eloranta

  • 12/01/97 The Volume Imaging Lidar has been moved to the Sheboygan site and installation of the beam steering unit is underway. We hope to begin system testing Tuesday afternoon. With luck, the lidar may be operational late on Wed. The telephone lines to the lidar van in Sheboygan are scheduled for installation Wed (12/02). The lidar voice line is 920-803-0605. If no answer at the lidar, try my office in Madison at 608-262-7327. Ed Eloranta

  • Return to the Index. Return to Lake-ICE
  • UW Lidar // Dec 2, 1997 // root@lidar.ssec.wisc.edu